# biblio.bib

### v1

@@ 1 Introduction
using the notation $\vec{X}^{N}:=(X^{i,N})_{1\leq i\leq N}$. The coefficients $b$ and $\sigma$ are globally Lipschitz functions which are path-dependent in a non-anticipative way, and $\sigma$ is globally bounded (precise conditions are listed in Assumption (A.1) below). As $N\rightarrow+\infty$, $\pi(\vec{X}^{N})$ converges to the path-law Ł$\L(\overline{X})$ of the McKean-Vlasov equation
$\mathrm{d}\overline{X}_{t}=b_{t}(\overline{X},\L(\overline{X}))\mathrm{d}t+\sigma_{t}(\overline{X},\L(\overline{X}))\mathrm{d}B^{1}_{t},\qquad X_{0}\sim X^{1}_{0}.$
@@
  • We compute the first-order fluctuations of $\pi(\vec{X}^{N})$ by proving a pathwise central limit theorem (CLT), i.e. the convergence of $N^{-1/2}[\pi(\vec{X}^{N})-\L(\overline{X})]$ towards a Gaussian field.
  • We prove a large deviation principle (LDP), estimating $\mathbb{P}(\pi(\vec{X}^{N})\in A)$ at the exponential scale for any measurable $A\subset\mathcal{P}(\mathcal{C}([0,T],\mathbb{R}^{d}))$, which may not contain Ł$\L(\overline{X})$.
@@ 1.2 Notations
  • $L^{p}_{\mathbb{P}}(\Omega,E)$ denotes the space of $E$-valued random variables $X:\Omega\rightarrow E$, on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ such that Ł$\L(X)$ belongs to $\mathcal{P}_{p}(E)$.
@@ 2 Statement of the main results
$\mathrm{d}\overline{X}_{t}=b_{t}(\overline{X},\L(\overline{X}))\mathrm{d}t+\sigma_{t}(\overline{X},\L(\overline{X}))\mathrm{d}B_{t}$
@@ 2.1 Discretised system with general driving noise
    If $W_{p}(\pi(\vec{\zeta}^{N},\vec{W}^{N}),\L(\zeta,{W}))\rightarrow 0$, $\mathbb{P}$-a.s. as $N\rightarrow+\infty$, then
    $W_{p}(\pi(\vec{X}^{h,N}),\L({X}^{h}))\rightarrow 0,\quad\mathbb{P}\text{-a.s. as }N\rightarrow+\infty.$
@@
    If the sequence of the Ł$\L(\pi(\vec{\zeta}^{N},\vec{W}^{N}))$ satisfies the LDP with good rate function $I:\mathcal{P}_{p}(\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})\rightarrow[0,+\infty]$, then the sequence of the Ł$\L(\pi(\vec{X}^{h,N}))$ satisfies the LDP with good rate function
@@ 2.2 Discretised setting with Brownian noise
$\mathrm{d}X^{h}_{t_{h}}=b_{t_{h}}(X^{h},\L(X^{h}))\mathrm{d}t+\sigma_{t_{h}}(X^{h},\L(X^{h}))\mathrm{d}B_{t},\qquad X^{h}_{0}=0,$
@@
$\mathrm{d}\delta X^{h}_{t_{h}}=\big{[}D_{x}\sigma_{t_{h}}(X^{h},\L(X^{h}))\cdot\delta X^{h}+\delta_{P}\sigma_{t_{h}}(X^{h},\L(X^{h}),\tilde{X}^{h})\big{]}\mathrm{d}B_{t}+\big{\{}D_{x}b_{t_{h}}(X^{h},\L(X^{h}))\cdot\delta X^{h}\\ \phantom{abcdefgh}+\delta_{P}b_{t_{h}}(X^{h},\L(X^{h}),\tilde{X}^{h})+\mathbb{E}\big{[}D_{y}\delta_{P}b_{t_{h}}(X^{h},\L(X^{h}),{X}^{h})\cdot\delta X^{h}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\big{\}}\mathrm{d}t\\ +\mathbb{E}\big{[}D_{y}\delta_{P}\sigma_{t_{h}}(X^{h},\L(X^{h}),{X}^{h})\cdot\delta X^{h}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}B_{t},\qquad\delta X^{h}_{0}=0.\phantom{abcd}$ (7)
@@ 2.3 Continuous setting with Brownian noise
$\mathrm{d}X_{t}=b_{t}(X,\L(X))\mathrm{d}t+\sigma_{t}(X,\L(X))\mathrm{d}B_{t},\qquad X_{0}=0,$
@@
$\mathrm{d}\delta X_{t}=\big{[}D_{x}\sigma_{t}(X,\L(X))\cdot\delta X+\delta_{P}\sigma_{t}(X,\L(X),\tilde{X})\big{]}\mathrm{d}B_{t}+\big{\{}D_{x}b_{t}(X,\L(X))\cdot\delta X\\ \phantom{abcdefgh}+\delta_{P}b_{t}(X,\L(X),\tilde{X})+\mathbb{E}\big{[}D_{y}\delta_{P}b_{t}(X,\L(X),X)\cdot\delta X\big{|}(\tilde{B}_{s})_{0\leq s\leq t}\big{]}\big{\}}\mathrm{d}t\\ +\mathbb{E}\big{[}D_{y}\delta_{P}\sigma_{t}(X,\L(X),X)\cdot\delta X\big{|}(\tilde{B}_{s})_{0\leq s\leq t}\big{]}\mathrm{d}B_{t},\qquad\delta X_{0}=0.\phantom{abcd}$ (8)
@@ 2.4 Contraction
is continuous, Proposition 2.7 and Theorem 2.9 induce by contraction the LDP for the sequence of the Ł$\L(\pi_{\cdot}(\vec{X}^{h,N}))$ with good rate function
@@
$\overline{I}_{0}(P_{\cdot})=H(P_{0}|\L(X^{1}_{0}))+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{1}{2}|v_{t}|^{2}\mathrm{d}P_{t}\mathrm{d}t,$
@@ 3.1 Construction of the discretised system
Moreover, for $(\zeta,W)\in L^{p}_{\mathbb{P}}(\Omega,\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})$, (4) has a unique solution in the sense of Definition 2.2, given by $\overline{S}^{h,\L(\zeta,W)}(\zeta,W)$.
In particular, $\overline{S}^{h,R}_{\#}R$ belongs to $\mathcal{P}_{p}(\mathcal{C}^{d})$. In the case $R=\L(\zeta,W)$, $\overline{S}^{h,R}_{\#}R$ is the law of the solution of (4). Following [Coghi2020PathwiseMT, Section 2], $\overline{S}^{h,R}_{\#}R$ can be characterised as the unique fixed-point of the map
@@
for $ih\leq t\leq(i+1)h$, uniquely extending the definition till time $(i+1)h$. All this makes sense because $b_{ih}(x,P)$ and $\sigma_{ih}(x,P)$ only depend on $x_{\wedge ih}$ and $P_{\wedge ih}$ using (A.1). By induction, the above Euler scheme builds $(X^{h}_{\wedge ih},\L(X^{h}_{\wedge ih}))$ for every $i\geq 0$, proving existence and pathwise uniqueness for (4). In particular, this proves existence and uniqueness for Ł$\L(X^{h})$.
To obtain an everywhere-defined solution, we set $\overline{S}^{h,R}(x_{0},\gamma):=S^{h,\mathcal{L}(X^{h})}(x_{0},\gamma)$, where Ł$\L(X^{h})$ is the path-law of the solution of (4) for any $(\zeta,W)$ with law $R$. $\overline{S}^{h,R}(\zeta,W)$ a.s. satisfies (9) with $P=\L(X^{h})$, so that $\overline{S}^{h,R}(\zeta,W)$ is the pathwise solution of (4). Consequently, $\overline{S}^{h,R}_{\#}R=\mathcal{L}(X^{h})$, and $\overline{S}^{h,R}(x_{0},\gamma)$ satisfies (11) for every $(x_{0},\gamma)$. Uniqueness stems from using Lemma 3.1-(i) with $P=\overline{S}^{h,R}_{\#}R$. ∎
@@
is continuous. Moreover, for $(\zeta,W)\in L^{p}_{\mathbb{P}}(\Omega,\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})$, $\Psi_{h}(\L(\zeta,W))$ is the law of the unique solution of (4).
@@
Using (A.1) and the bound (12), $\mathbb{E}_{\tilde{\mathbb{P}}}[\sup_{0\leq t\leq T}|b_{t}(X^{h,k},\L(X^{h,k}))|]$ is bounded uniformly in $k$. Taking expectations, we deduce that for every $\delta\in(0,h]$,
@@
Using this together with (12), [billingsley2013convergence, Theorem 7.3] now gives that the $\overline{S}^{h,R_{k}}_{\#}R_{k}=\L(X^{h,k})$ form a relatively compact sequence in $\mathcal{P}(\mathcal{C}^{d})$ for the weak topology.
Step 3. $W_{p}$-convergence. Since $R_{k}$ converges in the $W_{p}$-topology, $(|\zeta^{k}|^{p})_{k\geq 1}$ and $(|W^{k}|^{p})_{k\geq 1}$ are uniformly integrable [villani2009optimal, Definition 6.8-(iii)]. From (13), we then deduce that $(\sup_{0\leq t\leq T}|X^{h,k}_{t}|^{p})_{k\geq 1}$ is uniformly integrable. Using Step 2. and the characterisation [villani2009optimal, Definition 6.8-(iii)], this implies that (Ł$(\L(X^{h,k}))_{k\geq 1}$ is relatively compact in $(\mathcal{P}_{p}(\mathcal{C}^{d}),W_{p})$. Let $P$ be any limit point in $\mathcal{P}_{p}(\mathcal{C}^{d})$.
Using the $\tilde{\mathbb{P}}$-a.s. convergence of $(\zeta^{k},W^{k})$, and the continuity of $(x_{0},\gamma,P)\mapsto S^{h,P}(x_{0},\gamma)$, which is implied by Lemma 3.1-(ii), $\overline{S}^{h,R_{k}}(\zeta^{k},W^{k})=S^{h,\L(X^{h,k})}(\zeta^{k},W^{k})$ has a sub-sequence that $\tilde{\mathbb{P}}$-a.s. converges towards $S^{h,P}(\zeta,W)$. Along such a sub-sequence, we deduce that Ł$\L(X^{h,k})=\L(\overline{S}^{h,R_{k}}(\zeta^{k},W^{k}))$ weakly converges towards Ł$\L(S^{h,P}(\zeta,W))$, so that $P=S^{h,P}_{\#}R$. This implies that $S^{h,P}(\zeta,W)$ is a solution of (4). The pathwise uniqueness in Lemma 3.2 then gives that $S^{h,P}(\zeta,W)=\overline{S}^{h,R}(\zeta,W)$, $\tilde{\mathbb{P}}$-a.s., so that $P=\overline{S}^{h,R}_{\#}R=\Psi_{h}(R)$. The unique limit point in $W_{p}$ of the relatively compact sequence $(\overline{S}^{h,R_{k}}_{\#}R_{k})_{k\geq 1}$ is thus $\Psi_{h}(R)$; hence $\Psi_{h}(R_{k})=\overline{S}^{h,R_{k}}_{\#}R_{k}$ converges towards $\Psi_{h}(R)$. ∎
@@ 3.2 Mean-field limit and large deviations
${X}^{h}=S^{h,\L({X}^{h})}(\zeta,{W})\quad\text{and}\quad X^{h,i,N}=S^{h,\pi(\vec{X}^{h,N})}(\zeta^{i},W^{i}),$
@@
From the Sanov theorem in the $W_{p}$-topology [wang2010sanov, Theorem 1.1], the sequence of the Ł$\L(\pi(\vec{X}^{N}_{0},\vec{B}^{N}))$ satisfies the LDP with good rate function $R^{\prime}\in\mathcal{P}_{p}(\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})\mapsto H(R^{\prime}|R)$, where $R:=\L(X^{1}_{0},B^{1})$. By contraction [dupuis2011weak, Theorem 1.3.2], Corollary 2.5 proves that the Ł$\L(\pi(\vec{X}^{h,N}))$ satisfies the LDP with good rate function

### v2

@@ 1 Introduction
using the notation $\vec{X}^{N}:=(X^{i,N})_{1\leq i\leq N}$. The coefficients $b$ and $\sigma$ are globally Lipschitz functions which are path-dependent in a non-anticipative way, and $\sigma$ is globally bounded (precise conditions are listed in Assumption (A.1) below). As $N\rightarrow+\infty$, $\pi(\vec{X}^{N})$ converges to the path-law $\mathcal{L}(\overline{X})$ of the McKean-Vlasov equation
$\mathrm{d}\overline{X}_{t}=b_{t}(\overline{X},\mathcal{L}(\overline{X}))\mathrm{d}t+\sigma_{t}(\overline{X},\mathcal{L}(\overline{X}))\mathrm{d}B^{1}_{t},\qquad X_{0}\sim X^{1}_{0}.$
@@
  • We compute the first-order fluctuations of $\pi(\vec{X}^{N})$ by proving a pathwise central limit theorem (CLT), i.e. the convergence of $N^{-1/2}[\pi(\vec{X}^{N})-\mathcal{L}(\overline{X})]$ towards a Gaussian field.
  • We prove a large deviation principle (LDP), estimating $\mathbb{P}(\pi(\vec{X}^{N})\in A)$ at the exponential scale for any measurable $A\subset\mathcal{P}(\mathcal{C}([0,T],\mathbb{R}^{d}))$, which may not contain $\mathcal{L}(\overline{X})$.
@@ 1.2 Notations
  • $L^{p}_{\mathbb{P}}(\Omega,E)$ denotes the space of $E$-valued random variables $X:\Omega\rightarrow E$, on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ such that $\mathcal{L}(X)$ belongs to $\mathcal{P}_{p}(E)$.
@@ 2 Statement of the main results
$\mathrm{d}\overline{X}_{t}=b_{t}(\overline{X},\mathcal{L}(\overline{X}))\mathrm{d}t+\sigma_{t}(\overline{X},\mathcal{L}(\overline{X}))\mathrm{d}B_{t}$
@@ 2.1 Discretised system with general driving noise
    If $W_{p}(\pi(\vec{\zeta}^{N},\vec{W}^{N}),\mathcal{L}(\zeta,{W}))\rightarrow 0$, $\mathbb{P}$-a.s. as $N\rightarrow+\infty$, then
    $W_{p}(\pi(\vec{X}^{h,N}),\mathcal{L}({X}^{h}))\rightarrow 0,\quad\mathbb{P}\text{-a.s. as }N\rightarrow+\infty.$
@@
    If the sequence of the $\mathcal{L}(\pi(\vec{\zeta}^{N},\vec{W}^{N}))$ satisfies the LDP with good rate function $I:\mathcal{P}_{p}(\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})\rightarrow[0,+\infty]$, then the sequence of the $\mathcal{L}(\pi(\vec{X}^{h,N}))$ satisfies the LDP with good rate function
@@ 2.2 Discretised setting with Brownian noise
$\mathrm{d}X^{h}_{t_{h}}=b_{t_{h}}(X^{h},\mathcal{L}(X^{h}))\mathrm{d}t+\sigma_{t_{h}}(X^{h},\mathcal{L}(X^{h}))\mathrm{d}B_{t},\qquad X^{h}_{0}=0,$
@@
$\mathrm{d}\delta X^{h}_{t_{h}}=\big{[}D_{x}\sigma_{t_{h}}(X^{h},\mathcal{L}(X^{h}))\cdot\delta X^{h}+\delta_{P}\sigma_{t_{h}}(X^{h},\mathcal{L}(X^{h}),\tilde{X}^{h})\big{]}\mathrm{d}B_{t}+\big{\{}D_{x}b_{t_{h}}(X^{h},\mathcal{L}(X^{h}))\cdot\delta X^{h}\\ \phantom{abcdefgh}+\delta_{P}b_{t_{h}}(X^{h},\mathcal{L}(X^{h}),\tilde{X}^{h})+\mathbb{E}\big{[}D_{y}\delta_{P}b_{t_{h}}(X^{h},\mathcal{L}(X^{h}),{X}^{h})\cdot\delta X^{h}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\big{\}}\mathrm{d}t\\ +\mathbb{E}\big{[}D_{y}\delta_{P}\sigma_{t_{h}}(X^{h},\mathcal{L}(X^{h}),{X}^{h})\cdot\delta X^{h}\big{|}(\tilde{B}_{s})_{0\leq s\leq t_{h}}\big{]}\mathrm{d}B_{t},\qquad\delta X^{h}_{0}=0.\phantom{abcd}$ (7)
@@ 2.3 Continuous setting with Brownian noise
$\mathrm{d}X_{t}=b_{t}(X,\mathcal{L}(X))\mathrm{d}t+\sigma_{t}(X,\mathcal{L}(X))\mathrm{d}B_{t},\qquad X_{0}=0,$
@@
$\mathrm{d}\delta X_{t}=\big{[}D_{x}\sigma_{t}(X,\mathcal{L}(X))\cdot\delta X+\delta_{P}\sigma_{t}(X,\mathcal{L}(X),\tilde{X})\big{]}\mathrm{d}B_{t}+\big{\{}D_{x}b_{t}(X,\mathcal{L}(X))\cdot\delta X\\ \phantom{abcdefgh}+\delta_{P}b_{t}(X,\mathcal{L}(X),\tilde{X})+\mathbb{E}\big{[}D_{y}\delta_{P}b_{t}(X,\mathcal{L}(X),X)\cdot\delta X\big{|}(\tilde{B}_{s})_{0\leq s\leq t}\big{]}\big{\}}\mathrm{d}t\\ +\mathbb{E}\big{[}D_{y}\delta_{P}\sigma_{t}(X,\mathcal{L}(X),X)\cdot\delta X\big{|}(\tilde{B}_{s})_{0\leq s\leq t}\big{]}\mathrm{d}B_{t},\qquad\delta X_{0}=0.\phantom{abcd}$ (8)
@@ 2.4 Contraction
is continuous, Proposition 2.7 and Theorem 2.9 induce by contraction the LDP for the sequence of the $\mathcal{L}(\pi_{\cdot}(\vec{X}^{h,N}))$ with good rate function
@@
$\overline{I}_{0}(P_{\cdot})=H(P_{0}|\mathcal{L}(X^{1}_{0}))+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{1}{2}|v_{t}|^{2}\mathrm{d}P_{t}\mathrm{d}t,$
@@ 3.1 Construction of the discretised system
Moreover, for $(\zeta,W)\in L^{p}_{\mathbb{P}}(\Omega,\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})$, (4) has a unique solution in the sense of Definition 2.2, given by $\overline{S}^{h,\mathcal{L}(\zeta,W)}(\zeta,W)$.
In particular, $\overline{S}^{h,R}_{\#}R$ belongs to $\mathcal{P}_{p}(\mathcal{C}^{d})$. In the case $R=\mathcal{L}(\zeta,W)$, $\overline{S}^{h,R}_{\#}R$ is the law of the solution of (4). Following [Coghi2020PathwiseMT, Section 2], $\overline{S}^{h,R}_{\#}R$ can be characterised as the unique fixed-point of the map
@@
for $ih\leq t\leq(i+1)h$, uniquely extending the definition till time $(i+1)h$. All this makes sense because $b_{ih}(x,P)$ and $\sigma_{ih}(x,P)$ only depend on $x_{\wedge ih}$ and $P_{\wedge ih}$ using (A.1). By induction, the above Euler scheme builds $(X^{h}_{\wedge ih},\mathcal{L}(X^{h}_{\wedge ih}))$ for every $i\geq 0$, proving existence and pathwise uniqueness for (4). In particular, this proves existence and uniqueness for $\mathcal{L}(X^{h})$.
To obtain an everywhere-defined solution, we set $\overline{S}^{h,R}(x_{0},\gamma):=S^{h,\mathcal{L}(X^{h})}(x_{0},\gamma)$, where $\mathcal{L}(X^{h})$ is the path-law of the solution of (4) for any $(\zeta,W)$ with law $R$. $\overline{S}^{h,R}(\zeta,W)$ a.s. satisfies (9) with $P=\mathcal{L}(X^{h})$, so that $\overline{S}^{h,R}(\zeta,W)$ is the pathwise solution of (4). Consequently, $\overline{S}^{h,R}_{\#}R=\mathcal{L}(X^{h})$, and $\overline{S}^{h,R}(x_{0},\gamma)$ satisfies (11) for every $(x_{0},\gamma)$. Uniqueness stems from using Lemma 3.1-(i) with $P=\overline{S}^{h,R}_{\#}R$. ∎
@@
is continuous. Moreover, for $(\zeta,W)\in L^{p}_{\mathbb{P}}(\Omega,\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})$, $\Psi_{h}(\mathcal{L}(\zeta,W))$ is the law of the unique solution of (4).
@@
Using (A.1) and the bound (12), $\mathbb{E}_{\tilde{\mathbb{P}}}[\sup_{0\leq t\leq T}|b_{t}(X^{h,k},\mathcal{L}(X^{h,k}))|]$ is bounded uniformly in $k$. Taking expectations, we deduce that for every $\delta\in(0,h]$,
@@
Using this together with (12), [billingsley2013convergence, Theorem 7.3] now gives that the $\overline{S}^{h,R_{k}}_{\#}R_{k}=\mathcal{L}(X^{h,k})$ form a relatively compact sequence in $\mathcal{P}(\mathcal{C}^{d})$ for the weak topology.
Step 3. $W_{p}$-convergence. Since $R_{k}$ converges in the $W_{p}$-topology, $(|\zeta^{k}|^{p})_{k\geq 1}$ and $(|W^{k}|^{p})_{k\geq 1}$ are uniformly integrable [villani2009optimal, Definition 6.8-(iii)]. From (13), we then deduce that $(\sup_{0\leq t\leq T}|X^{h,k}_{t}|^{p})_{k\geq 1}$ is uniformly integrable. Using Step 2. and the characterisation [villani2009optimal, Definition 6.8-(iii)], this implies that $(\mathcal{L}(X^{h,k}))_{k\geq 1}$ is relatively compact in $(\mathcal{P}_{p}(\mathcal{C}^{d}),W_{p})$. Let $P$ be any limit point in $\mathcal{P}_{p}(\mathcal{C}^{d})$.
Using the $\tilde{\mathbb{P}}$-a.s. convergence of $(\zeta^{k},W^{k})$, and the continuity of $(x_{0},\gamma,P)\mapsto S^{h,P}(x_{0},\gamma)$, which is implied by Lemma 3.1-(ii), $\overline{S}^{h,R_{k}}(\zeta^{k},W^{k})=S^{h,\mathcal{L}(X^{h,k})}(\zeta^{k},W^{k})$ has a sub-sequence that $\tilde{\mathbb{P}}$-a.s. converges towards $S^{h,P}(\zeta,W)$. Along such a sub-sequence, we deduce that $\mathcal{L}(X^{h,k})=\mathcal{L}(\overline{S}^{h,R_{k}}(\zeta^{k},W^{k}))$ weakly converges towards $\mathcal{L}(S^{h,P}(\zeta,W))$, so that $P=S^{h,P}_{\#}R$. This implies that $S^{h,P}(\zeta,W)$ is a solution of (4). The pathwise uniqueness in Lemma 3.2 then gives that $S^{h,P}(\zeta,W)=\overline{S}^{h,R}(\zeta,W)$, $\tilde{\mathbb{P}}$-a.s., so that $P=\overline{S}^{h,R}_{\#}R=\Psi_{h}(R)$. The unique limit point in $W_{p}$ of the relatively compact sequence $(\overline{S}^{h,R_{k}}_{\#}R_{k})_{k\geq 1}$ is thus $\Psi_{h}(R)$; hence $\Psi_{h}(R_{k})=\overline{S}^{h,R_{k}}_{\#}R_{k}$ converges towards $\Psi_{h}(R)$. ∎
@@ 3.2 Mean-field limit and large deviations
${X}^{h}=S^{h,\mathcal{L}({X}^{h})}(\zeta,{W})\quad\text{and}\quad X^{h,i,N}=S^{h,\pi(\vec{X}^{h,N})}(\zeta^{i},W^{i}),$
@@
From the Sanov theorem in the $W_{p}$-topology [wang2010sanov, Theorem 1.1], the sequence of the $\mathcal{L}(\pi(\vec{X}^{N}_{0},\vec{B}^{N}))$ satisfies the LDP with good rate function $R^{\prime}\in\mathcal{P}_{p}(\mathbb{R}^{d}\times\mathcal{C}^{d^{\prime}})\mapsto H(R^{\prime}|R)$, where $R:=\mathcal{L}(X^{1}_{0},B^{1})$. By contraction [dupuis2011weak, Theorem 1.3.2], Corollary 2.5 proves that the $\mathcal{L}(\pi(\vec{X}^{h,N}))$ satisfies the LDP with good rate function